​(Bond valuation)  Calculate the value of a bond that matures in 17 years and has a $1,000 par value. The annual coupon interest rate is 13 percent and the​ market's required yield to maturity on a​ comparable-risk bond is 15 percent.

The value of the bond is $___. (Round to the nearest​ cent.)
Answer #1

Annual coupon=1000*13%=130

Hence value of bond=Annual coupon*Present value of annuity factor(15%,17)+$1000*Present value of discounting factor(15%,17)

=130*6.047160755+$1000*0.092925886

=$879.06(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=130[1-(1.15)^-13]/0.15

=130*6.047160755

2.Present value of discounting factor=1000/1.15^17

=1000*0.092925886